DIMETHYL ETHER THREE DIMENTIONAL SPECTRA

\underline{MARIA VILLA}, Departamento de Quimica. Universidad Autonoma Metropolitana, Av. San Rafael Atlixco 186, Col Vicentina, Iztapala, Mexico D.F. 09340, MEXICO; MA. LUISA SENENT, Departamento de Astrofisica Molecular e Infrarroja, Instituto de Estructura de la Materia, C.S.I.C., Serrano 113B, Madrid 28006, SPAIN.

The dimethyl ether is a non-rigid molecule with three internal motions: the two methyl rotations and the COC bending. This motions are studied using MP2 and CCSD(T) calculations using a aug-cc-pVTZ basis set and applying the ZPVE corrections. With these results the 3D spectra is calculated.